# RIFT: Closed-Loop RL Fine-Tuning for Realistic and Controllable Traffic Simulation

### v2

@@ B.1 Experiment Framework
Our framework for reliable AV-centric closed-loop simulation is developed upon well-established traffic simulation platforms, notably the CARLA Leaderboard CARLA_Leaderboard and Bench2Drive jia2024bench2drive , which serve as standard benchmarks in autonomous driving research. Traditionally, these platforms use predefined scenarios along the AV’s global route to evaluate the multi-dimensional performance of AV methods. In contrast, we replace these static scenarios with dynamically generated traffic flows by randomly spawning background vehicles around the AV’s global path and simulating their behavior using rule-based driving policies, as described in Section˜3.1. Through the CBV identification mechanism outlined in Section˜B.2, we naturally introduce interactions between the AV and CBVs, thereby generating continuous, interactive scenarios over time. This framework serves as the foundation for both the training and evaluation processes in this paper.
@@ B.2 Route-level Analysis for CBV Identification
where, $\hat{D}_{\text{global}}(p_{k},\mathcal{V}_{i})$ and $\hat{D}_{\text{global}}(p_{k},\mathcal{V}_{\text{AV}})$ denote the estimated travel distance required for the background vehicle $\mathcal{V}_{i}$ and the AV to reach waypoint $p_{k}$, respectively. The distance-to-goal for each vehicle is computed by determining the distance from its current position to the target waypoint $p_{k}$ using the A* global path planning algorithm hart1968a_star . A threshold $\delta$ is introduced to define the maximum allowable difference in distance-to-goal. A background vehicle is considered critical and included in the CBV set $\mathcal{C}$ if the absolute distance-to-goal difference between it and the AV is smaller than $\delta$.
@@ B.3 Algorithm Framework
For clarity, we summarize the procedure of RIFT within our AV-centric closed-loop simulation framework in Algorithm 1. The planning model initializes from the IL pre-trained checkpoint provided by Pluto official codebase [^1] https://github.com/jchengai/pluto, followed by closed-loop RL fine-tuning within the CARLA simulator carla to generate realistic and controllable traffic scenarios.
@@
1:Input: IL pre-trained planning model $\pi_{\theta_{\text{init}}}$, buffer $\mathcal{D}$ $\triangleright$ IL pre-training (nuPlan caesar2021nuplan )
@@
3:for $\text{iteraction}=1,\dots,I$ do $\triangleright$ Closed-loop RL fine-tuning (CARLA carla )
@@
Generate $G$ trajectories $\{\tau_{i}\}_{i=1}^{G}\sim\pi_{\theta_{old}}$ for each CBV $\triangleright$ Policy inference
@@
Update model $\pi_{\theta}$ by maximizing the RIFT objective (Equation˜8)
@@ B.4 Training Details
We perform closed-loop fine-tuning on selected modules of the IL pre-trained planning model (Pluto). As shown in the ablation results (Section˜5.4), fine-tuning only the trajectory scoring head achieves the best trade-off between realism and controllability. Accordingly, all fine-tuning baselines adopt this setting to ensure consistency and fair comparison. Our training framework is built on the open-source Lightning platform [^2] https://github.com/Lightning-AI/pytorch-lightning. Fine-tuning is conducted on $2\times$ Bench2Drive220, while evaluation is performed on dev10, both from the Bench2Drive project. All experiments are conducted on NVIDIA GeForce RTX 4090D GPUs, with each fine-tuning run taking approximately 8 hours on a single GPU. Detailed training setups and hyperparameter configurations are provided in Table˜6 and Table˜6.
@@ B.5 Baselines Detailed Description
  • Pluto cheng2024pluto is an open-source IL-based planning framework for autonomous driving. It processes vectorized scene representations as input and outputs multimodal trajectories for downstream planning. In the AV-centric closed-loop simulation, the method directly uses a pre-trained checkpoint without additional fine-tuning.
  • FREA chen2024frea is an RL-based approach designed to generate safety-critical yet AV-feasible scenarios. It incorporates a feasibility-aware training objective. In the AV-centric closed-loop simulation, FREA selects potential collision points along the AV’s global route as adversarial goals.
  • PPO chen2024frea is a variant of FREA that focuses solely on generating safety-critical scenarios. Unlike FREA, it disregards the feasibility constraints of AV and treats adversariality as the only optimization objective.
  • FPPO-RS chen2024frea is another FREA variant that integrates AV’s feasibility constraints into the reward shaping process, thereby balancing adversariality with scenario reasonability.
  • PPO-Pluto fine-tunes the pre-trained planning model using the PPO algorithm PPO . The fine-tuning follows the same reward structure as detailed in Section˜B.7, aligning with RIFT.
  • REINFORCE-Pluto employs the REINFORCE algorithm sutton1999reinforce to fine-tune the pre-trained Pluto model under the same reward design as detailed in Section˜B.7.
  • GRPO-Pluto utilizes the basic GRPO algorithm shao2024grpo for fine-tuning, employing the pre-trained Pluto model as the reference for KL regularization, while incorporating the standard PPO-Clip.
  • SFT-Pluto is a purely supervised fine-tuning approach, where PDM-Lite Beibwenger2024pdmLite serves as the expert model, providing supervision at the target speed level.
  • RTR-Pluto zhang2023rtr is a hybrid framework combining imitation and reinforcement learning. While the original RTR utilizes human driving trajectories as supervision, our setting replaces this with PDM-Lite due to the lack of human-level demonstrations. The RL component uses sparse infraction-based rewards, consistent with the original RTR, and applies PPO for optimization.
  • RS-Pluto peng2024improving also adopts a hybrid IL+RL paradigm, originally trained via REINFORCE using ground-truth supervision and sparse rewards to ensure safety and realism. In our adaptation, PDM-Lite substitutes the ground-truth expert, while the rest of the methodology remains unchanged.
@@ B.6 Forward Simulation
Trajectory-based imitation learning often neglects underlying system dynamics, leading to discrepancies between planned and executed behavior cheng2024planTF . To address this, we perform a closed-loop forward simulation for each candidate trajectory ${\tau}_{i}$, yielding a rollout $\widetilde{\tau}_{i}$. The simulation integrates a PID controller for trajectory tracking with a kinematic bicycle model for state propagation. The PID controller is identical to that used during closed-loop execution, ensuring behavioral consistency between training and deployment, while the simulation time step is aligned with the temporal resolution of candidate trajectories. By evaluating rollouts rather than raw candidate trajectories, we mitigate this gap and obtain more reliable assessments.
@@ B.7 State-Wise Reward Model Setup
To capture diverse human driving styles, we decompose driving behaviors into distinct reward components, following cusumano2025robust . Different styles are constructed by combining weights assigned to each reward component (detailed in Table˜6), enabling a range of behaviors from aggressive to conservative. The total driving reward is defined as:
@@
$r^{t}_{i}=\mathrm{StateWiseRM}(\phi(\widetilde{\tau}^{t}_{i})).$ (22)
All features, except the infraction indicators, are directly derived from the rollouts. To estimate future infractions, we follow the forecasting model in Beibwenger2024pdmLite to simulate other agents’ future positions based on current states and actions and identify collisions via bounding box overlap. Off-road violations are detected using the Euclidean Signed Distance Field (ESDF) cheng2024pluto .
@@ B.8 Controllability and Realism Metrics
Kinematic Metrics. Following montali2023wosac , kinematic realism is typically evaluated against ground-truth trajectories. As CARLA provides no expert demonstrations, we adopt distribution-level metrics chen2024review ; huang2025IntSim to assess CBV behavior in terms of speed and acceleration. Specifically, we employ three measures—the Shapiro–Wilk test on speed (S-SW), the Wasserstein Distance on speed (S-WD), and the Shapiro–Wilk test on acceleration (A-SW)—defined as follows:
  • Wasserstein Distance (WD) vaserstein1969wd : measures the distance between two distributions $\mu$ and $\nu$. Since CARLA provides a predefined target speed for agents, we use WD to compare the simulated CBV speed distribution with the target speed distribution as the reference.
@@
  • Shapiro–Wilk test (SW)shapiro1965sw : evaluates the normality of speed and acceleration distributions—a simplifying assumption supported by empirical traffic studieszhang2023rtr ; yan2023NeuralNDE —to capture the statistical naturalness of CBV motion.
@@
Interaction Metrics. Following the metric design principles proposed in the WOSAC challenge montali2023wosac and other widely adopted evaluation frameworks lin2024ccdiff ; chen2024review , we adopt a set of well-established metrics to comprehensively evaluate agent interactions:
  • Collision Per Kilometer (CPK) chen2024review : the average number of scenario collisions per kilometer of driving distance.
  • Route Progress (RP) chen2024review : the total distance traveled by all CBVs, reflecting route completion.
  • 2D Time-to-Collision (2D-TTC) guo20232dttc : the minimum of longitudinal and lateral time-to-collision from the AV’s perspective, capturing the immediate interaction risk posed by CBVs.
  • Anticipated Collision Time (ACT) venthuruthiyil2022act : a safety-critical metric measuring the AV’s proximity to potential collisions, reflecting the interaction intensity introduced by CBVs.
@@
  • Off-Road Rate (ORR) chen2024review : the percentage of time that CBVs spend off-road on average.
@@ C.1 AV Methods Implementation
  • PDM-Lite Beibwenger2024pdmLite : A rule-based privileged expert method that achieves state-of-the-art performance on the CARLA Leaderboard 2.0 by leveraging components such as the Intelligent Driver Model and the kinematic bicycle model. This open-source method serves as a strong baseline for comparison.
  • PlanT Renz2022PlanT : An explainable, learning-based planning method that operates on an object-level input representation and is trained through imitation learning.
  • UniAD hu2023_uniad : A planning-oriented unified framework integrating perception, prediction, mapping, and planning into one end-to-end model using query-based interfaces.
  • VAD jiang2023vad : A fast, end-to-end vectorized driving paradigm representing scenes with vectorized motion and map elements for efficient, safe planning.
@@ C.2 AV Evaluation Metrics
As detailed in Sections˜B.1 and B.4, we develop an AV-centric closed-loop simulation environment, including a training and evaluation pipeline based on Bench2Drive. The AV closed-loop evaluation metrics proposed in Bench2Drive extend the original metrics of the CARLA Leaderboard by emphasizing the specific strengths and weaknesses of different methods across various aspects, such as merging and overtaking, thereby making them suitable for evaluating performance under predefined scenarios. However, as noted in Section˜B.1, replacing predefined scenarios with CBV-generated traffic scenarios precludes the evaluation of specific AV capabilities. To systematically assess the quality of traffic scenarios generated by different CBV methods, we follow the practice of KING king and introduce PDM-Lite Beibwenger2024pdmLite —a rule-based privileged planner—as a reference AV. By measuring its performance under various CBV methods, we evaluate:
@@
These metrics enable a principled comparison of traffic quality generated by different CBV methods. Furthermore, to assess the capacity of generated traffic scenarios to expose AV limitations, we test multiple learning-based AV methods under an identical CBV method and quantify their relative performance drop compared to PDM-Lite Beibwenger2024pdmLite . The relative driving score degradation ($\Delta{DS}$) reflects how effectively the traffic scenario stresses the AV policy, with larger drops indicating stronger capability in revealing planning weaknesses.
@@ D.1 Detailed Qualitative Results of Style-Level Controllability
Quantitative results in Table˜3 show that the aggressive variant achieves greater driving efficiency at the expense of more frequent collisions and off-road events. To complement these findings, Figure˜5 presents a qualitative comparison in a single-lane intersection scenario where a leading BV halts at a stop sign. The aggressive CBV variant attempts an overtaking maneuver, resulting in a collision, whereas the normal CBV variant yields and waits, demonstrating distinct behavioral patterns induced by different reward preferences. These results highlight the controllability of RIFT in modulating driving style according to user-specified reward configuration.

### v3

@@ B.1 Experiment Framework
Our framework for reliable AV-centric closed-loop simulation is developed upon well-established traffic simulation platforms, notably the CARLA Leaderboard [71] and Bench2Drive [49], which serve as standard benchmarks in autonomous driving research. Traditionally, these platforms use predefined scenarios along the AV’s global route to evaluate the multi-dimensional performance of AV methods. In contrast, we replace these static scenarios with dynamically generated traffic flows by randomly spawning background vehicles around the AV’s global path and simulating their behavior using rule-based driving policies, as described in Section˜3.1. Through the CBV identification mechanism outlined in Section˜B.2, we naturally introduce interactions between the AV and CBVs, thereby generating continuous, interactive scenarios over time. This framework serves as the foundation for both the training and evaluation processes in this paper.
@@ B.2 Route-level Analysis for CBV Identification
where, $\hat{D}_{\text{global}}(p_{k},\mathcal{V}_{i})$ and $\hat{D}_{\text{global}}(p_{k},\mathcal{V}_{\text{AV}})$ denote the estimated travel distance required for the background vehicle $\mathcal{V}_{i}$ and the AV to reach waypoint $p_{k}$, respectively. The distance-to-goal for each vehicle is computed by determining the distance from its current position to the target waypoint $p_{k}$ using the A* global path planning algorithm [72]. A threshold $\delta$ is introduced to define the maximum allowable difference in distance-to-goal. A background vehicle is considered critical and included in the CBV set $\mathcal{C}$ if the absolute distance-to-goal difference between it and the AV is smaller than $\delta$.
@@ B.3 Algorithm Framework
For clarity, we summarize the procedure of RIFT within our AV-centric closed-loop simulation framework in Algorithm˜1. The planning model is initialized from the IL pre-trained checkpoint provided by Pluto official codebase [^1] https://github.com/jchengai/pluto, followed by RL fine-tuning within the CARLA simulator [64] to generate realistic and controllable traffic scenarios.
@@
1:Input: IL pre-trained planning model $\pi_{\theta_{\text{init}}}$, buffer $\mathcal{D}$ $\triangleright$ IL pre-training (nuPlan [15])
@@
3:for $\text{iteraction}=1,\dots,I$ do $\triangleright$ RL fine-tuning (CARLA [64])
@@
Obtain $G$ candidate trajectories $\{\tau_{i}\}_{i=1}^{G}$ from $\pi_{\theta_{old}}$ for each CBV $\triangleright$ Policy inference
@@
Update model $\pi_{\theta}$ by maximizing the RIFT objective (Equation˜9)
@@ B.4 Training Details
We perform RL fine-tuning on selected modules of the IL pre-trained planning model (Pluto). As shown in the ablation results (Section˜5.4), fine-tuning only the trajectory scoring head achieves the best trade-off between realism and controllability. Accordingly, all fine-tuning baselines adopt this setting to ensure consistency and fair comparison. Our training framework is built on the open-source Lightning platform [^2] https://github.com/Lightning-AI/pytorch-lightning. Fine-tuning is conducted on $2\times$ Bench2Drive220, while evaluation is performed on dev10, both from the Bench2Drive project. All experiments are conducted on NVIDIA GeForce RTX 4090D GPUs, with each fine-tuning run taking approximately 8 hours on a single GPU. Detailed training setups and hyperparameter configurations are provided in Table˜6 and Table˜6.
@@ B.5 Baselines Detailed Description
  • Pluto [56] is an open-source IL-based planning framework for autonomous driving. It processes vectorized scene representations as input and outputs multimodal trajectories for downstream planning. In the AV-centric closed-loop simulation, the method directly uses a pre-trained checkpoint without additional fine-tuning.
  • FREA [18] is an RL-based approach designed to generate safety-critical yet AV-feasible scenarios. It incorporates a feasibility-aware training objective. In the AV-centric closed-loop simulation, FREA selects potential collision points along the AV’s global route as adversarial goals.
  • PPO [18] is a variant of FREA that focuses solely on generating safety-critical scenarios. Unlike FREA, it disregards the feasibility constraints of AV and treats adversariality as the only optimization objective.
  • FPPO-RS [18] is another FREA variant that integrates AV’s feasibility constraints into the reward shaping process, thereby balancing adversariality with scenario reasonability.
  • PPO-Pluto fine-tunes the pre-trained planning model using the PPO algorithm [60]. The fine-tuning follows the same reward structure as detailed in Section˜B.7, aligning with RIFT.
  • REINFORCE-Pluto employs the REINFORCE algorithm [66] to fine-tune the pre-trained Pluto model under the same reward design as detailed in Section˜B.7.
  • GRPO-Pluto utilizes the basic GRPO algorithm [24] for fine-tuning, employing the pre-trained Pluto model as the reference for KL regularization, while incorporating the standard PPO-Clip.
  • SFT-Pluto is a purely supervised fine-tuning approach, where PDM-Lite [65] serves as the expert model, providing supervision at the target speed level.
  • RTR-Pluto [25] is a hybrid framework combining imitation and reinforcement learning. While the original RTR utilizes human driving trajectories as supervision, our setting replaces this with PDM-Lite due to the lack of human-level demonstrations. The RL component uses sparse infraction-based rewards, consistent with the original RTR, and applies PPO for optimization.
  • RS-Pluto [26] also adopts a hybrid IL+RL paradigm, originally trained via REINFORCE using ground-truth supervision and sparse rewards to ensure safety and realism. In our adaptation, PDM-Lite substitutes the ground-truth expert, while the rest of the methodology remains unchanged.
@@ B.6 Forward Simulation
Trajectory-based imitation learning often overlooks underlying system dynamics, leading to discrepancies between planned and executed behavior [73]. To address this issue, we perform a forward simulation for each candidate trajectory ${\tau}_{i}$ of the CBV, yielding a rollout $\widetilde{\tau}_{i}$. The simulation couples a PID controller for trajectory tracking with a kinematic bicycle model for state propagation. The PID controller is identical to that used during closed-loop execution, ensuring behavioral consistency between training and deployment. By evaluating rollouts rather than raw trajectories, we reduce this dynamics gap and obtain more reliable assessments.
In parallel, we also forecast the motions of surrounding actors. During data collection, the current actions $a^{\text{bg}}$ of surrounding actors are recorded. Following the rule-based forecasting scheme in [65], these actions are assumed constant over the forecast horizon and are used to advance surrounding states. The resulting actor forecasts are combined with the CBV rollouts to compute rewards, thereby ensuring that interaction effects with the environment are faithfully captured in evaluation.
While subsequent rollout is open-loop, the first transition is closed-loop. This step integrates (i) the same PID policy as in real execution, (ii) the observed current actions of surrounding actors, and (iii) a kinematic bicycle model that approximates CARLA’s single-step dynamics. Accordingly, the transition from $(s,a,a^{\text{bg}})$ to $s^{\prime}$ produces a reward consistent with the standard RL structure $(s,a)\to s^{\prime}\to r$. Subsequent rollout steps serve as open-loop estimates of longer-horizon outcomes, enriching evaluation while preserving closed-loop fidelity at the transition boundary.
@@ B.7 State-Wise Reward Model Setup
To capture diverse human driving styles, we decompose driving behaviors into distinct reward components, following [74]. Different styles are constructed by combining weights assigned to each reward component (detailed in Table˜6), enabling a range of behaviors from aggressive to conservative. The total driving reward is defined as:
@@
$r^{t}_{i}=\mathrm{StateWiseRM}(\phi(\widetilde{\tau}^{t}_{i}),s).$ (19)
All features, except the infraction indicators, are directly derived from the rollouts. To estimate future infractions, we follow the forecasting model in [65] to simulate other agents’ future positions based on current states and actions and identify collisions via bounding box overlap. Off-road violations are detected by projecting the rollout trajectory onto the HD map and checking its occupancy relative to the drivable area polygon set.
@@ B.8 Controllability and Realism Metrics
Kinematic Metrics. Following [42], kinematic realism is typically evaluated against ground-truth trajectories. As CARLA provides no expert demonstrations, we adopt distribution-level metrics [67, 76] to assess CBV behavior in terms of speed and acceleration. Specifically, we employ three measures—the Shapiro–Wilk test on speed (S-SW), the Wasserstein Distance on speed (S-WD), and the Shapiro–Wilk test on acceleration (A-SW)—defined as follows:
  • Wasserstein Distance (WD) [77]: measures the distance between two distributions $\mu$ and $\nu$. Since CARLA provides a predefined target speed for agents, we use WD to compare the simulated CBV speed distribution with the target speed distribution as the reference.
@@
  • Shapiro–Wilk test (SW) [78]: evaluates the normality of speed and acceleration distributions—a simplifying assumption supported by empirical traffic studies [25, 79]—to capture the statistical naturalness of CBV motion.
@@
Interaction Metrics. Following the metric design principles proposed in the WOSAC challenge [42] and other widely adopted evaluation frameworks [45, 67], we adopt a set of well-established metrics to comprehensively evaluate agent interactions:
  • Collision Per Kilometer (CPK) [67]: the average number of scenario collisions per kilometer of driving distance.
  • Route Progress (RP) [67]: the total distance traveled by all CBVs, reflecting route completion.
  • 2D Time-to-Collision (2D-TTC) [68]: the minimum of longitudinal and lateral time-to-collision from the AV’s perspective, capturing the interaction risk posed by CBVs.
  • Anticipated Collision Time (ACT) [69] : a safety-critical metric measuring the AV’s proximity to potential collisions, reflecting the interaction intensity introduced by CBVs.
@@
  • Off-Road Rate (ORR) [67]: the percentage of time CBVs spend off-road on average.
@@ C.1 AV Methods Implementation
  • PDM-Lite [65]: A rule-based privileged expert method that achieves state-of-the-art performance on the CARLA Leaderboard 2.0 by leveraging components such as the Intelligent Driver Model and the kinematic bicycle model. This open-source method serves as a strong baseline for comparison.
  • PlanT [70]: An explainable, learning-based planning method that operates on an object-level input representation and is trained through imitation learning.
  • UniAD [53]: A planning‑oriented unified framework integrating perception, prediction, mapping, and planning into one end‑to‑end model using query-based interfaces.
  • VAD [54]: A fast, end‑to‑end vectorized driving paradigm representing scenes with vectorized motion and map elements for efficient, safe planning.
@@ C.2 AV Evaluation Metrics
As detailed in Sections˜B.1 and B.4, we develop an AV-centric closed-loop simulation environment, including a training and evaluation pipeline based on Bench2Drive. The AV closed-loop evaluation metrics proposed in Bench2Drive extend the original metrics of the CARLA Leaderboard by emphasizing the specific strengths and weaknesses of different methods across various aspects, such as merging and overtaking, thereby making them suitable for evaluating performance under predefined scenarios. However, as noted in Section˜B.1, replacing predefined scenarios with CBV-generated traffic scenarios precludes the evaluation of specific AV capabilities. To systematically assess the quality of traffic scenarios generated by different CBV methods, we follow the practice of KING [17] and introduce PDM-Lite [65]—a rule-based privileged planner—as a reference AV. By measuring its performance under various CBV methods, we evaluate:
@@
These metrics enable a principled comparison of traffic quality generated by different CBV methods. Furthermore, to assess the capacity of generated traffic scenarios to expose AV limitations, we test multiple learning-based AV methods under an identical CBV method and quantify their relative performance drop compared to PDM-Lite [65]. The relative driving score degradation ($\Delta{DS}$) reflects how effectively the traffic scenario stresses the AV policy, with larger drops indicating stronger capability in revealing planning weaknesses.
@@
C.3 End-to-End AV Visualization
To further validate the feasibility of RIFT as a closed-loop evaluation framework, we extend its application beyond traffic scenario generation to testing end-to-end autonomous driving algorithms. In contrast to conventional adversarial approaches that often introduce unrealistic or overly aggressive behaviors, RIFT generates traffic flows that preserve the realism of human driving styles, engage the autonomous vehicle in genuine interactive behaviors, and maintain feasibility by ensuring that the resulting scenes, though diverse and stress-inducing, remain solvable for the end-to-end method. This balance enables RIFT to evaluate robustness under credible conditions while avoiding degenerate or unsolvable scenarios.
Figure˜5 presents representative interactions between RIFT-generated traffic flows and two representative end-to-end driving models, UniAD and VAD. As shown, RIFT adapts seamlessly to different driving policies, producing realistic and interactive scenes where the autonomous vehicle must negotiate with surrounding traffic. These results underscore the capability of RIFT to provide realistic yet interactive closed-loop evaluations, highlighting its potential as a versatile tool for testing the robustness of end-to-end AV systems.
Figure 5: Representative closed-loop interactions between RIFT-generated traffic flows and end-to-end autonomous driving algorithms. UniAD (top) and VAD (bottom) are shown interacting with surrounding vehicles orchestrated by RIFT, which preserves realistic driving styles while enabling dynamic CBV–AV interactions. The controlled background vehicle (CBV) is highlighted in purple, the autonomous vehicle (AV, end-to-end) in red, and other background vehicles (BVs) in blue.
@@ D.1 Detailed Qualitative Results of Style-Level Controllability
Quantitative results in Table˜3 show that the aggressive variant achieves greater driving efficiency at the expense of more frequent collisions and off-road events. To complement these findings, Figure˜6 presents a qualitative comparison in a single-lane intersection scenario where a leading BV halts at a stop sign. The aggressive CBV variant attempts an overtaking maneuver, resulting in a collision, whereas the normal CBV variant yields and waits, demonstrating distinct behavioral patterns induced by different reward preferences. These results highlight the controllability of RIFT in modulating driving style according to user-specified reward configuration.